Calculate the pH of a solution prepared by dissolving 0.876 moles of carbonic acid (H2CO3) and 0.543 moles of sodium hydrogen carbonate (NaHCO3) in water sufficient to yield 1.00 L of solution. The Ka of carbonic acid is 4.5x10^-7. Explain answer please
Answer #1

no of moles of carbonic acid used = 0.876 moles

no of moles of sodium hydrogen carbonate used = 0.543 moles

reaction between carbonic acid and sodium hydrogen carbonate is given bellow

H2CO3 + 2NaHCO3   Na2CO3 + 2H2O + 2CO2  

according to reaction 2 moles of sodium hydrogen carbonate reacts with 1 mole of carbonic acid

therefore 0.543 moles of sodium hydrogen carbonate will react with = 0.543/2 = 0.2715 mole H2CO3

but 0.876 moles of carbonic acid are taken which are excess than required

no of moles of carbonic acid remained unreacted = 0.876 - 0.543 = 0.333 moles

total volume of solution = 1liter

molarity = no of moles/volume in liter , substitute values to find molarity of carbonic acid

molarity of carbonic acid = 0.333/1 = 0.333M

carbonic acid dissociates as

H2CO3  + H2O HCO3-+ H3O+

Ka = [HCO3-][H3O+] / [H2CO3]

but  [HCO3-] = [H3O+] = x

Ka = [x][x] / [H2CO3] , given Ka of H2CO3 = 4.5 x 10-7  

Substitute the value in equation

4.5 10-7 = [x]2/ 0.333

[x]2 = 4.5 10-7 0.333 = 1.4985 x 10-7

[x] = 3871.0463 10-7

Concentration of H3O+ = 3871.0463 10-7 M

pH = - log[H+]

pH = - log (3871.0463 10-7) = 3.41

ans pH of resulatant solution formed is 3.41